OpenStudy (anonymous):

I need help what is 4 1/8 divided by 5 2/5 equivalent to

OpenStudy (anonymous):

turn it into a decimal. it helps when you use a calculator.

OpenStudy (anonymous):

4 1/8 = 4.125 5 2/5 = 5.4 Dividing them gives you: 0.76388888888 0.76388888888 = 0.76 0.76 = 76/100 = 19/25

OpenStudy (anonymous):

@Amyissweet medal if helped :D

OpenStudy (midhun.madhu1987):